OpenStudy (anonymous):

Simplify: (5x2 + 3x + 4) + (5x2 + 5x − 1)

OpenStudy (goldphenoix):

Combine like terms. You can't combine anything in the parenthesis, so you take out the parenthesis. It should look like: 5x^2 + 3x + 4 + 5x^2 + 5x -1. Now like I said, combine like terms. :) Do you know how to do that?

OpenStudy (anonymous):

No.

jimthompson5910 (jim_thompson5910):

something like 3x and 5x are like terms because they have the same variable with the same exponent so adding them means you just add the numbers in front and tack on an x when you're done so, 3x + 5x = 8x (notice how 3+5 = 8)

jimthompson5910 (jim_thompson5910):

other examples 6x + 4x = 10x 8x + 21x = 29x 13x^2 + 7x^2 = 20x^2 3xy + 9xy = 12xy

jimthompson5910 (jim_thompson5910):

hopefully all that makes sense

OpenStudy (anonymous):

Thank you guys. (:

jimthompson5910 (jim_thompson5910):

you're welcome, what did you get for the final answer